# We Have a Package for You! A Comprehensive Analysis of Package Hallucinations by Code Generating LLMs

### v1

@@ Abstract
The reliance of popular programming languages such as Python and JavaScript on centralized package repositories and open-source software, combined with the emergence of code-generating Large Language Models (LLMs), has created a new type of threat to the software supply chain: package hallucinations. These hallucinations, which arise from fact-conflicting errors when generating code using LLMs, represent a novel form of package confusion attack that poses a critical threat to the integrity of the software supply chain. This paper conducts a rigorous and comprehensive evaluation of package hallucinations across different programming languages, settings, and parameters, exploring how different configurations of LLMs affect the likelihood of generating erroneous package recommendations and identifying the root causes of this phenomena. Using 16 different popular code generation models, across two programming languages and two unique prompt datasets, we collect 576,000 code samples which we analyze for package hallucinations. Our findings reveal that 19.7% of generated packages across all the tested LLMs are hallucinated, including a staggering 205,474 unique examples of hallucinated package names, further underscoring the severity and pervasiveness of this threat. We also implemented and evaluated mitigation strategies based on Retrieval Augmented Generation (RAG), self-detected feedback, and supervised fine-tuning. These techniques demonstrably reduced package hallucinations, with hallucination rates for one model dropping below 3%. While the mitigation efforts were effective in reducing hallucination rates, our study reveals that package hallucinations are a systemic and persistent phenomenon that pose a significant challenge for code generating LLMs.
@@ 1 Introduction
Modern generative AI models are large deep learning systems, pre-trained on extensive datasets, enabling them to learn the underlying distribution and produce novel outputs. Today, a variety of commercial and open-source generative models are available, capable of producing synthetic images (e.g., Stable Diffusion [70], DALL-E [68], MidJourney [54]), videos (e.g., OpenAI Sora [63], Microsoft VASA-1 [87]), audio (e.g., AudioGen [33]) and text/conversations (e.g., ChatGPT [5], BlenderBot [72]) through easy-to-use interfaces and textual commands. These powerful models have enabled a range of new applications such as creative content generation, personalized virtual assistants, sophisticated chatbots, and coding assistants.
@@
A particular class of generative models, known as Large Language Models (or LLMs), are foundational text generation systems capable of understanding and creating natural language and other types of textual content, and can be trained to perform a wide range of specialized tasks such as answering questions, summarizing documents, translating languages, and completing sentences. Some popular examples of such foundational LLMs include GPT-4 [1], Gemini [76], LlaMA [80], and Mistral [29]. One such emergent use of LLMs is the ability to generate computer code, accomplished by training or fine-tuning on vast programming-related datasets including code repositories, technical forums, coding platforms, documentation, and web data relevant to programming. Both commercial/black-box (e.g., GPT-4 [1], Claude [3]) and open-source (e.g., CodeLlama [71], DeepSeek Coder [19]) varieties of such code generation LLMs are readily available, and are being extensively used by both novice and expert programmers in their coding workflows to increase productivity. Studies indicate that up to 97% of developers are using generative AI to some degree and that approximately 30% of code written today is AI-generated, reflecting significant perceived gains in efficiency and convenience [73, 43].
@@
Despite their tremendous success in solving complex language-related tasks, LLMs have many shortcomings, with a phenomenon called hallucinations being a particularly serious issue. Hallucinations are outputs produced by LLMs that may appear to be plausible or truthful, but in reality are either factually incorrect, overestimate and distort the true facts, are nonsensical, or completely unrelated to the input task. Hallucinations present a critical challenge to the effective and safe deployment of LLMs in public facing applications due to their potential to generate inaccurate or misleading information. As a result, there has been increased attention in the research literature on the topics of detection and mitigation of hallucinations in LLMs [22, 28]. However, most of the existing research efforts have focused only on hallucinations in classical natural language generation and prediction tasks such as machine translation, summarization, and conversational AI [24, 53, 40, 10]. The occurrence and impact of hallucinations during code generation, particularly regarding the type of hallucinated content and its implications for code security, remain relatively unexplored in the research literature. Hallucinations occurring during LLM-assisted code generation could result in code snippets that conflict with the user’s requirements, the contextual information provided at input time, or the code knowledge base itself. Such hallucinations and their effects, if not appropriately mitigated during generation or time-of-use, could seriously undermine the correctness, security, and performance of the developed software. Recently, Liu et al. [46] has shown that popular LLMs (e.g., ChatGPT, CodeRL and CodeGen) significantly hallucinate during code generation and have established a comprehensive taxonomy of hallucinations in LLM-generated code.
@@
In this work, we are concerned with a very specific type of hallucination during code generation called package hallucination. Package hallucination occurs when an LLM generates code that either recommends or contains a reference to a package/library which does not actually exist. Package hallucinations, especially if they are repeated, can have significant implications on the security of the generated code, with one severe and emergent threat being a package confusion attack. In a package confusion attack, an adversary takes advantage of package hallucinations in the LLM-generated code by actually publishing a package with the same name as the hallucinated or fictitious package and containing some malicious code/functionality. Then, as other unsuspecting and trusting LLM users are subsequently recommended the same fictitious package in their generated code, they end up downloading the adversary-created malicious package, resulting in a successful compromise. From that point on, any other software or code that utilizes or is dependent on this code containing the hallucinated package, is automatically also dependent on the adversary-created malicious package, thus becoming part of an entire codebase or software dependency chain.
@@
Package confusion attacks, through techniques such as typosquatting (i.e., creating packages with names similar to popular ones to deceive users) and name similarity, have been a longstanding issue in the open-source software community. Package hallucinations by code-generating LLMs threaten to exacerbate the problem by exposing an additional threat surface for such attacks [58, 74, 34]. Trivial cross-referencing based hallucinated package detection techniques (i.e., comparing the package name with a list of known legitimate packages) are practically infeasible due to the large, dynamic, and open-source nature of popular package repositories. Additionally, an adversary may have already created a package corresponding to a hallucinated package making cross-referencing ineffective. Despite recent research [46] showing that LLMs are prone to package hallucinations, the extent to which this phenomenon occurs in current state-of-the-art (SOTA) publicly-available commercial (black-box) and open-source LLMs, the nature of these hallucinations, their impact on the security of the generated code, and the effectiveness of potential mitigation measures have not been comprehensively investigated.
@@
In this paper we conduct the first systematic study of the frequency and nature of package hallucinations across a variety of code generation LLMs operating under a diverse set of model settings and parameters. We rigorously analyze and quantify the impact of these hallucinations on the security of the generated code (vis-à-vis package confusion attacks) and evaluate several mitigation strategies. We specifically make the following novel contributions:
@@
  • Quantifying the incidence and origins of package hallucinations in code generating LLMs: We comprehensively analyze the prevalence of package hallucinations in Python and JavaScript code generated by popular publicly-available commercial and open-source LLMs. We also examine the common behaviors in LLMs that lead to package hallucinations, including hallucination repetition, output verbosity, and the ability of models to detect their own hallucinations.
@@
  • Analyze the effect of model settings and training data on package hallucinations: We further study how specific model settings, such as training data recency, model temperature, and decoding strategies affect the occurrence and nature of package hallucinations.
@@
  • Characterizing the generated hallucinated packages: We carefully study several key properties of the hallucinated packages, such as their semantic similarity to popular packages, the occurrence of package recommendations from other programming languages, package persistence, and the significance of packages that have been recently removed from source repositories.
@@
  • Mitigation Strategies: We propose and comprehensively evaluate several techniques to effectively mitigate package hallucinations in LLM-generated code.
@@ 2 Background and Related Work
In this section we provide a brief background on code generation LLMs, the issue of hallucinations in LLMs, and its impact on open-source software security.
2.1 Large Language Models (LLMs)
LLMs and Automated Code Generation using LLMs. The advent of extremely versatile and high-performing attention-based transformer models [81] have catalyzed the emergence of foundational language models, which are large-scale models described by billions of parameters and pre-trained on vast datasets. These foundational LLMs can be tailored to perform a wide range of downstream tasks without the need for task-specific architecture modifications [4]. Models such as BERT (Bidirectional Encoder Representations from Transformers) [11], CLIP (Contrastive Language-Image Pretraining) [67], and GPT (Generative Pre-Trained Transformer) [5] are examples of foundational language models that have emerged as new benchmarks for a wide range of tasks.
Hallucinations by LLMs. It has been well-documented that LLMs can unintentionally produce harmful information [48, 85], be manipulated for malicious purposes [20, 31], expose private information [41], and often carry inherent biases in their training data [15]. A related phenomenon is the notion of hallucinations, which are instances where LLMs generate misleading or entirely fictitious information. These errors arise in various forms: the model might misinterpret the intended input (input-conflicting hallucination), produce inconsistencies with prior outputs (context-conflicting hallucination), or contradict established facts (fact-conflicting hallucination) [89]. Hallucinations can stem from three main root causes: (i) data, (ii) training, and (iii) inference [23]. Hallucinations from data occur when the data source itself is flawed with misinformation [44], bias [15], or incomplete records [61]. Architecture flaws [45] or sub-optimal training objectives [83] during training could also result in downstream hallucinations, while inference time issues such as defective coding strategies [21] and imperfect decoding representations [8, 49] also contribute to hallucinations.
The persistent issue of hallucinations in LLMs has spurred extensive research into various mitigation strategies, broadly categorized into prompt engineering and model architecture enhancements [79]. Prompt engineering techniques such as Retrieval Augmented Generation (RAG) [39] and self-refinement methods [52] aim to refine the input provided to the model to produce more accurate outputs. Alternatively, developing more robust models involves approaches such as supervised fine-tuning [77], inference time intervention [42], and incorporating knowledge graphs [27] to improve model understanding and reduce errors/hallucinations in the model output. The manipulation of the temperature parameter within each model has also been shown to significantly influence the prevalence of hallucinations in LLM output [2]. The temperature parameter modulates the probability distribution over potential output tokens [55]. Lower temperatures produce less random, more predictable outputs, while higher temperatures increase the likelihood of sampling low-frequency tokens, raising the risk of hallucinations [12].
The conflict between the non-deterministic nature of LLMs and mitigating hallucinations is inherently challenging. This non-determinism fosters creativity and generates diverse, innovative content, making it a desired feature despite its role in producing hallucinations. Balancing creativity with accuracy remains a central challenge in the deployment of LLMs, underscoring the complexity of developing effective mitigation strategies.
2.2 Software Supply Chain Security
Open-Source Software (OSS) and Software Supply Chain. OSS is any software or application with source code that anyone can inspect, modify, and enhance. Reliance on OSS, in particular publicly-available packages and libraries for software development, has been steadily increasing. The Linux foundation estimates that 98% of the current codebases include some form of OSS [57], and this dependence is only expected to increase during the next year [69]. Many programming languages, including four of the top-5 most popular on GitHub, rely on centralized package repositories to manage package distribution and dependency management. Services such as npm [59] and PyPI [13] are the two most popular examples of package repositories, serving as the primary package ecosystem for Python and JavaScript, respectively.
OSS Security. The convenience and simplicity afforded by the public nature of open-source software repositories also makes them an ideal platform for malware distribution by malicious actors. A total of 245,000 malicious packages were found on open-source repositories in 2023, a 300% increase from 2022 and 2×2$\times$2 × the amount of all previous years combined [73], another trend that is predicted to continue for the foreseeable future [16]. Existing OSS security vulnerabilities can be broadly categorized as either insecure code or malicious code. Insecure code is the result of poor coding practices or unmaintained legacy code that allows malicious actors to exploit an open vulnerability. Repository statistics show that one in eight open-source packages have security risks and 18% of available packages are unmaintained [73]. On the other hand, malicious packages or libraries contain intentionally harmful functionality, which research has shown allows attackers to execute arbitrary code on victim machines once these packages are downloaded [60].
Package Confusion Attacks. Many packages/libraries rely on other packages to function, thus creating huge dependency trees which further compounds the problem of OSS security [90]. For an adversary to infect a particular software product or ecosystem, they only need to infect a single package in the dependency chain [58, 32]. Code generating LLMs only further exacerbate this problem by generating insecure code, which is typically not scrutinized by users due to the high level of trust in the underlying model [65]. Further compounding the issue is the potential inclusion of this insecure code in the dependency chain of other packages, leading to a cascading effect where vulnerabilities are propagated across numerous codebases. There have been many well-documented OSS supply chain attacks on both public and private package repositories and dependency chains, including the notable SolarWinds compromise [6]. Public OSS repositories such as PyPI and npm have implemented several measures to raise the barrier for malicious packages to be distributed on their platforms, such as two-factor authentication, namespace protection, and software signing [82, 88]. It is currently not known if the repositories utilize scanning methods for malicious code, but they often do not release the full list of packages that have been removed.
@@ 2.2 Software Supply Chain Security
Once a malicious package has been created and uploaded to a software repository, adversaries have devised various techniques to trick users into downloading the package, which is subsequently incorporated into their codebase and dependency chains. At a high level, these attacks work by deliberately naming malicious packages to mimic or appear as legitimate packages, and are collectively referred to as package confusion attacks [58]. Package confusion attacks can be broadly categorized into typosquatting, combosquatting, brandjacking, and similarity attacks [35], and are distinct from other types of software supply chain attacks such as corrupting legitimate packages or developing unique malicious packages from scratch as part of a long-term campaign. There have been over 1,200 documented package confusion attacks in the last six years [58], including an incident involving the popular PyTorch package [74] and a campaign by the North Korean Advanced Persistent Threat Lazarus Group [34]. At one point the PyPI repository had to temporarily suspend package uploads and new account creation due to the number of malicious packages being uploaded [13].
@@
Code-generating LLMs and the hallucinated packages recommended by them significantly amplify these existing challenges and risks associated with package confusion attacks.
@@ 2.3 Related Work
The idea of code generation models recommending malicious or typosquatted packages was first suggested in 2021 as tools such as GPT-3 and Codex were released as viable code generation platforms [5]. At the time, the risk of code generation tools suggesting vulnerable, malicious, or typosquatted packages was assessed to be low [9]. Notably, the related but distinct concept of package hallucinations was not explicitly considered in this initial risk assessment; either because such an attack had not been considered or the threat was thought to be negligible. The capabilities of generative AI agents have advanced significantly since that introductory evaluation. Although a thorough study of package hallucinations has not been previously conducted, a blog post by Lanyado in June 2023 [37], and subsequent follow-up post in March 2024 [36], present the first preliminary findings on the threat of package hallucination attacks. Their first study reports that package hallucinations occur in 20% of JavaScript responses and 35% of Python responses using ChatGPT 3.5. In the follow-up work, they found that 24.2% of package recommendations were hallucinated using GPT-4 and 22.2% using GPT-3.5. Two other commercial models were tested, Google Gemini and Cohere, which reported 64.5% and 29.1% hallucination rates, respectively. No open-source models were tested. This study also contained a proof of concept for executing a package hallucination attack by actually publishing a package (with dummy code) using the name of a hallucinated package that was generated during their testing. In just three months, the dummy package received over 30,000 downloads on PyPI, establishing that the attack vector exposed by LLM-generated package hallucinations is absolutely viable [36].
Compared to these previous works which primarily focused on identifying the prevalence of hallucinations and providing high-level recommendations, our work aims to significantly advance the understanding of package hallucinations in LLM-generated code. We conduct a rigorous and comprehensive evaluation across a broader range of models (including the first analysis of package hallucination in open-source models of any kind) at a scale and level of detail that has not been previously documented. To this end, we provide thorough testing with a larger custom dataset covering two programming languages (Python and JavaScript) followed by analysis into the root causes and defining characteristics of this phenomenon, including the first study on the effects of temperature, decoding strategies, deleted packages, stale training data, semantic similarity, model verbosity, and language confusion on package hallucinations. We then propose potential mitigation strategies to curb hallucination attacks, the first such work with recommendations to limit hallucinations in code generation LLMs. Lastly, we contribute a novel dataset for code generation that can be used to test LLMs across a broad range of coding tasks.

### v2

@@ Abstract
The reliance of popular programming languages such as Python and JavaScript on centralized package repositories and open-source software, combined with the emergence of code-generating Large Language Models (LLMs), has created a new type of threat to the software supply chain: package hallucinations. These hallucinations, which arise from fact-conflicting errors when generating code using LLMs, represent a novel form of package confusion attack that poses a critical threat to the integrity of the software supply chain. This paper conducts a rigorous and comprehensive evaluation of package hallucinations across different programming languages, settings, and parameters, exploring how a diverse set of models and configurations affect the likelihood of generating erroneous package recommendations and identifying the root causes of this phenomenon. Using 16 popular LLMs for code generation and two unique prompt datasets, we generate 576,000 code samples in two programming languages that we analyze for package hallucinations. Our findings reveal that that the average percentage of hallucinated packages is at least 5.2% for commercial models and 21.7% for open-source models, including a staggering 205,474 unique examples of hallucinated package names, further underscoring the severity and pervasiveness of this threat. To overcome this problem, we implement several hallucination mitigation strategies and show that they are able to significantly reduce the number of package hallucinations while maintaining code quality. Our experiments and findings highlight package hallucinations as a persistent and systemic phenomenon while using state-of-the-art LLMs for code generation, and a significant challenge which deserves the research community’s urgent attention.
@@ 1 Introduction
Modern generative AI models, pre-trained on extensive multi-modal datasets, can create synthetic images (e.g., Stable Diffusion [65], DALL-E [64], MidJourney [52]), videos (e.g., OpenAI Sora [60], Microsoft VASA-1 [81]), audio (e.g., AudioGen [31]) and text/conversations (e.g., ChatGPT [5], BlenderBot [67]). These models have enabled a range of new applications such as creative content generation, personalized virtual assistants, and sophisticated chatbots. Foundational Large Language Models (LLMs), including GPT-4 [1] and LlaMA [74], are trained on enormous and diverse textual datasets, and typically excel in a variety of natural language tasks, such as answering questions, summarizing documents, and translating languages.
@@
An emerging application of these LLMs is to generate computer code, which is typically accomplished by first training or fine-tuning them using vast amounts of programming/code datasets found on online repositories (e.g., GitHub), technical forums, and documentation. Both commercial/black-box (e.g., GPT-4 [1], Claude [3]) and open-source (e.g., CodeLlama [66], DeepSeek Coder [17]) varieties of such code-generating LLMs are readily available and being extensively used by both novice and expert programmers in their coding workflows to increase productivity. Recent studies indicate that up to 97% of developers are using generative AI to some degree and that approximately 30% of code written today is AI-generated, reflecting significant perceived gains in efficiency and convenience [68, 41].
@@
One serious shortcoming of LLMs is a phenomenon referred to as hallucination. Hallucinations are outputs produced by LLMs that are factually incorrect, nonsensical, or completely unrelated to the input task. Hallucinations present a critical obstacle to the effective and safe deployment of LLMs in public-facing applications due to their potential to generate inaccurate or misleading information. As a result, there has been increased efforts to research the detection and mitigation of hallucinations in LLMs [20, 26]. However, most existing research has focused only on hallucinations in classical natural language generation and prediction tasks such as machine translation, summarization, and conversational AI [22, 51, 38, 9]. The occurrence and impact of hallucinations during code generation, particularly regarding the type of hallucinated content and its implications for code security, are still in the nascent stages of research. Recently, Liu et al. [44] have shown that popular LLMs (e.g., ChatGPT, CodeRL, and CodeGen) significantly hallucinate during code generation and have established a taxonomy of hallucinations in LLM-generated code.
@@
In this work, we focus on a specific type of hallucination during code generation called package hallucination. Package hallucination occurs when an LLM generates code that recommends or contains a reference to a package that does not actually exist. An adversary can exploit package hallucinations, especially if they are repeated, by publishing a package to an open-source repository with the same name as the hallucinated or fictitious package and containing some malicious code/functionality. Then, as other unsuspecting and trusting LLM users are subsequently recommended the same fictitious package in their generated code, they end up downloading the adversary-created malicious package, resulting in a successful compromise. This compromise can then spread through an entire codebase or software dependency chain, infecting any code that relies on the malicious package. This is a variation of the classical package confusion attack that has been enabled by code-generating LLMs.
@@
Package confusion attacks, through techniques such as typosquatting (i.e., creating packages with names similar to popular ones to deceive users) and name similarity, have been a longstanding issue in the open-source software community [55, 69, 32]. Package hallucinations by code-generating LLMs threaten to exacerbate the problem by exposing an additional threat surface for such attacks. Trivial cross-referencing methods (i.e., comparing a generated package name with a list of known packages) are ineffective for detecting a package hallucination attack, as an adversary may have already published the hallucinated package with malicious code. Open source repositories make no guarantee about the safety of hosted content; the mere presence of a package in an open-source repository does not confirm its credibility. A recent blog post [34] suggests that LLMs are prone to package hallucinations and provides a first approximation of its prevalence, but the extent to which this phenomenon occurs in state-of-the-art (SOTA) commercial and open-source LLMs, the nature of these hallucinations, and the effectiveness of potential mitigation measures have not been comprehensively investigated before.
@@
In this paper we conduct the first systematic study of the frequency and nature of package hallucinations across a variety of code generation LLMs operating under a diverse set of model settings and parameters. We specifically make the following novel contributions:
@@
  • Quantifying the frequency and prevalence of package hallucinations by code generating LLMs: We first comprehensively analyze the prevalence of package hallucinations in Python and JavaScript code generated by popular publicly-available commercial and open-source LLMs. We also examine the common behaviors in LLMs that lead to package hallucinations, including hallucination repetition, output verbosity, and the ability of models to detect their own hallucinations.
@@
  • Analyzing the effect of model settings and training data on package hallucinations: We further study how specific model settings, such as training data recency, model temperature, and decoding strategies affect the occurrence and nature of package hallucinations.
@@
  • Characterizing the common traits of the generated hallucinated packages: We carefully study several key properties of the hallucinated packages, such as their semantic similarity to popular packages, the occurrence of package recommendations from other programming languages, package persistence, and the influence of recently removed packages on hallucination rate.
@@
  • Mitigation Strategies: We propose and comprehensively evaluate several techniques to effectively mitigate package hallucinations in LLM-generated code while maintaining the ability to produce effective code, something that has not been done before.
  • Data We make publicly available a novel dataset of 19,500 coding prompts for a vast range of coding tasks, 586,000 generated coding samples for fine-tuning/analysis, and a list of 205,474 hallucinated package names that can be used to monitor for potential instances of a package hallucination attack.
@@ 2 Background and Related Work
In this section we provide a brief background on open-source software security, code generating LLMs, and the issue of hallucinations in LLMs.
2.1 Background
Package Confusion Attacks in Open-source Software Repositories. Modern software development has seen an increased reliance on open-source software packages and libraries available on public centralized repositories. Many modern programming languages now rely on such centralized package repositories, with PyPI [12] (for Python) and npm [56] (for JavaScript) being the two most popular repositories. The open nature of these repositories, where anyone can upload new code packages/libraries, makes them an attractive platform for malware distribution. For instance, a total of 245,000 malicious code packages were discovered on open-source software repositories in 2023 alone [68].
@@ 2.1 Background
Once a malicious package is uploaded, adversaries employ various techniques to trick users into downloading it, thereby integrating it into their codebases and dependency chains. These attacks often involve deliberately naming malicious packages to mimic legitimate ones, a tactic known as a package confusion attack [55]. Package confusion attacks can be broadly categorized into typosquatting, combosquatting, brandjacking, and similarity attacks [33], and are distinct from other types of software supply chain attacks such as corrupting legitimate packages or developing unique malicious packages from scratch as part of a long-term campaign. More than 1,200 package confusion attacks have been documented in the last six years [55], including the notable PyTorch compromise [69] and the Lazarus Group campaign [32].
Packages/libraries often rely on other packages to function, thus creating extensive dependency trees, infecting a single package in this dependency chain can be sufficient to compromise an entire software product or ecosystem [55, 30]. Public OSS repositories such as PyPI and npm have implemented various measures, including two-factor authentication, namespace protection, and software signing to mitigate the distribution of malicious packages [76, 82]. However, it remains unclear whether these repositories utilize any scan-based techniques for detecting malicious code, and they often do not disclose the full list of removed packages.
Automated Code Generation using LLMs. Modern LLMs are continuing to demonstrate advanced source code generation capabilities, with success rates for correctly answering coding prompts surging from 25% in June 2021 to 96% by April 2024 [8]. With the increasing use of these models for software development, concerns are rising about the likelihood of producing insecure or incorrect code that could create vulnerabilities in deployed applications. Early versions of code generation LLMs were found to generate code containing vulnerabilities listed in the MITRE Top-25 Common Weakness Enumeration (CWE) 40% of the time [61]. Moreover, recent research has shown that AI-assisted programming not only results in less secure code but also instills a false sense of security among developers [62].
Hallucinations by LLMs. It has been well-documented that LLMs can unintentionally produce harmful information [46, 79], be manipulated for malicious purposes [18, 29], expose private information [39], and carry inherent biases in their training data [14]. A related phenomenon is hallucinations, where LLMs generate misleading or entirely fictitious information. These errors take various forms: the model might misinterpret the intended input (input-conflicting hallucination), produce inconsistencies with prior outputs (context-conflicting hallucination), or contradict established facts (fact-conflicting hallucination) [83]. Hallucinations can arise from three main root causes: (i) data, (ii) training, and (iii) inference [21]. Data-related hallucinations occur when the source data itself is flawed with misinformation [42], bias [14], or incomplete records [58]. Architecture flaws [43] or suboptimal training objectives [77] during training could also result in downstream hallucinations, while inference time issues such as defective coding strategies [19] and imperfect decoding representations [7, 47] are further contributors. The probabilistic nature of LLMs presents a challenge in mitigating hallucinations. This nondeterminism, while it fosters creativity and generates diverse and innovative content, also contributes to the generation of hallucinations. Balancing creativity with accuracy remains a central challenge in deploying LLMs, underscoring the complexity of developing effective mitigation strategies.
Package Hallucinations and Security Risks. Package hallucinations, a special form of fact-conflicting hallucinations, are instances where LLMs generate fictitious (non-existent) or erroneous package names in the generated source code. As outlined earlier, an adversary can quickly create malicious packages (on the appropriate open-source repository) with the same name as these hallucinated packages, thus effecting a very simple, yet effective, form of package confusion attack. Unsuspecting users, who trust the LLM output, may not scrutinize the validity of these hallucinated packages in the generated code and could inadvertently include these malicious packages in their codebase [62]. This resulting insecure open-source code also has the potential of being included in the dependency chain of other packages and code, leading to a cascading effect where vulnerabilities are propagated across numerous codebases. The simplicity and scale of such LLM-enabled package confusion attacks highlight the critical need for quantifying this existing risk, understanding the nature of this unique type of hallucination, and developing effective mitigation techniques that maintain the utility of the code generated by the LLMs. This is precisely what we aim to accomplish in this work.
@@ 2.2 Related Work
The idea of code generation models recommending malicious or typosquatted packages was first suggested in 2021 as tools such as GPT-3 and Codex were released as viable code generation platforms [5]. At the time, the risk of code generation tools suggesting vulnerable, malicious, or typosquatted packages was assessed to be low [8]. In particular, the related but distinct concept of package hallucinations was not explicitly considered in this initial risk assessment; either because such an attack had not been considered or because the threat was thought to be negligible. The capabilities of generative AI agents have advanced significantly since that introductory evaluation.
Although a comprehensive study on the prevalence of package hallucinations in LLM-generated code has not been done previously, a recent blog post by Lanyado [35] outlines the results of some preliminary tests on commercial LLMs such as GPT, Gemini, and Cohere. Their testing does confirm the presence of hallucinated packages in the code generated by these LLMs but their initial estimate of hallucination rate is 5x higher than our findings and they do not consider popular open-source LLMs or study possible mitigation approaches. They also fail to systematically characterize the generated hallucinated packages and model properties that have an impact on hallucinations.
In contrast, we conduct a rigorous and comprehensive evaluation across a broader range of models, including the first analysis of package hallucinations in open-source models of any kind, at a scale that has not been done previously. To this end, we provide thorough testing with a larger custom dataset covering two programming languages (namely, Python and Javascript), followed by a detailed analysis of the significant characteristics of this phenomenon.